OpenStudy (anonymous):

If you wanted to make the graph of y = 6x + 7 less steep, which equation could you use?

OpenStudy (anonymous):

A. y = 2x + 7 B. y = 6x + 4 C. y = -6x + 7 D. y = 10x + 7

OpenStudy (anonymous):

those are the choices..

OpenStudy (mathstudent55):

The larger the absolute value of the slope is, the steeper the line is.